OpenStudy (anonymous):

At the beginning of the year, Addison Company's assets are $273,000 and its equity is $204,750. During the year, assets increase $80,000 and liabilities increase $42,000. What is the equity at the end of the year?

OpenStudy (sandra):

The formula for "equity" is assets - liabilities?

OpenStudy (anonymous):

Yes.

OpenStudy (ash2326):

Assets- liability= equity Let's find the liability at the beginning of the year 273000-liability=204750 liabilities=$68250 Now assets increase 80,000 so Present assets =273000+80000=$351000 Liabilities increased by 42000 Present Liabilities = 68250+42000= $110250 Equity=Assets-Liabilities Equity=351000-110250= $240750

OpenStudy (anonymous):

Hope you get the help Tania08

OpenStudy (anonymous):

I hope so too!

OpenStudy (anonymous):

So the answer is $240750